OpenStudy (anonymous):

(6^x)/(6^x-2) = 8...I know you have to use logs, but I'm not sure how

OpenStudy (anonymous):

6^x=8(6^x-2) 7*6^x=16 6^x=16/7 xlog6=log(16/7) x=log(16/7)/log(6)=0.461378096

OpenStudy (anonymous):

I don't get the second step

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

6^x=8(6^x-2) 6^x=8*6^x-16 8*6^x-6^x=16 7*6^x=16

OpenStudy (anonymous):

ok i got it thank you